The free-fall acceleration on the moon is 1.62 m/s2 . What is the length of a pendulum whose period on the moon matches the period of a 1.70-m-long pendulum on the earth? Lmoon= _____ m
Answer #1


formua for time period of simple pendulum is given by

T = 2pi sqrt(L/g)

for length 1.7 m

time period on earth is

T = 2* 3.14* sqrt(1.70/9.81)

T = 2.614 secs

now for this period on moon, the length Corresponds to

T = 2pi * sqrt(L/g)

sqrt(L/g) = 2.614/(2pi)

L/g = 0.416^2

L = 0.173056 * 1.62

L = 0.28 m or 28 cm --------<<<<<<<<<<Answer